# Learning to Think: Information-Theoretic Reinforcement Fine-Tuning for LLMs

### v1

@@ 6 Conclusion
This supplementary material provides results for additional experiments and details to reproduce our results that could not be included in the paper submission due to space limitations.
@@ Appendix A Proofs
Proof. For a given state $s_{k}$, the posterior distribution $p(\tilde{\theta}_{k}\mid s_{k})\propto p(s_{k}\mid\tilde{\theta}_{k})\;p(\tilde{\theta}_{k})$, take the logarithm to get the unstandardized posterior $L(\tilde{\theta}_{k})=\log p(s_{k}\mid\tilde{\theta}_{k})+\log p(\tilde{\theta}_{k})$. Among them,
@@ Appendix B Pseudo-Code
Input: Task distribution $P_{task}$; base model $\pi_{\theta}$;
@@
1:  for iteration = $1,\cdots,T$ do
2:     Sample a batch of tasks $X\sim P_{task}$
3:     for each $x\in X$ do
4:        $s_{1}\leftarrow\text{prompt}(x)$
5:        for episodes = $2,\cdots,K$ do
6:           Sample $N$ rollouts $\{z_{k}^{i},y_{k}^{i}\}\}_{i=1}^{N}\sim\pi_{\theta}(\cdot|s_{k})$
7:           Compute average progress reward $r_{k}^{\rm prg}$ via Definition 4.1 and Theorem 4.2
8:           Compute average outcome reward $r_{k}^{\rm out}$ following [16]
9:        end for
10:     end for
11:     Update the policy $\pi_{\theta}$ via Eq.5
12:  end for
13:  return solution
@@ Appendix D Benchmark Datasets
In this section, we briefly introduce all datasets used in our experiments. In summary, the benchmark datasets can be divided into two categories: (i) reasoning tasks for mathematical derivation, including AIME24–25, AMC, MATH500 [18], MinervaMATH [27], and Omni-MATH [13]; and (ii) reasoning tasks for code generation via HumanEval [8] The compositions of these benchmarks are as follows:
  • AIME24–25 comprises 30 questions from the 2024 and 2025 American Invitational Mathematics Examination, with 15 fill-in-the-blank questions per exam. These questions are more difficult than AMC, spanning number theory, combinatorics, geometry, and algebra.
@@
  • MinervaMATH comprises 12,500 high-school–level contest questions. Each includes detailed solution steps and spans prealgebra through precalculus.
@@ Appendix E Implementation Details
The implementation of our approach is built upon the openRL codebase and Llama-Factory codebase, with custom modifications to the loss function, i.e., replacing the reward function to our proposed reward mentioned in Subsection 4.2. For model training, we directly load the base models from Hugging Face, including DeepScaleR-1.5B-Preview, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeekR1-Distill-Qwen-7B, and Qwen2-7B-Instruct. For different reasoning tasks, we introduce the experimental settings in the corresponding sections of Section 5 and Appendix F. Unless otherwise specified, we follow the protocol of each benchmark and record the maj@4 results across different models. The training configuration is: the learning rate is set to $1.0\times 10^{-6}$, with a cosine learning rate scheduler, and a warm-up ratio of 0.1. We use a batch size of 256, with a maximum prompt length of 4,096 tokens and a maximum completion length of 16,384 tokens. The model is trained for 1 epoch, up to 10 epochs. Additionally, we set the ‘use_vllm’ flag to True to enable vLLM acceleration, with a GPU memory utilization of 0.8. We also utilize mixed precision training with BF16 enabled. The parameters for compression penalty approximation are handled by a single-layer MLP, with a Fisher information matrix damping factor set to $10^{-5}$. The regularization hyperparameters $\alpha$ and $\beta$ are set to 0.8 and 0.6 according to grid search results, respectively. Also, $\alpha$ can be set to 1 for simplicity, with the performance drop less than 1%. More evaluation of implementation configurations is provided in Appendix F. The entire training is conducted on A100 GPU clusters, ensuring scalability and high computational efficiency. The main code is provided in supplementary materials.
@@ Appendix F Additional Experiments and Full Results
+outcome-reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (–1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
+length penalty 40.3 (–2.5) 30.3 (–6.4) 77.3 (–5.7) 83.2 (-2.0) 23.0 (-1.6) 50.8 (–3.7)
@@
Figure 9: Ablation study on batch size. We evaluate batch sizes of 64, 128, 256.
@@ F.1 More Details and Results of the Motivating Experiments
In Subsection 3.2, we evaluate how efficiently existing methods use tokens. We benchmark two base models, DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B, on Omni-MATH (4,428 questions across 33+ subfields, split into Tiers 1–4 by expert difficulty labels). Both models have been fine-tuned with outcome-based RL. To study performance at varying reasoning depths, we split each generated reasoning chain into up to $K=20$ episodes using ‘<think>…</think>’, then record the sequential-generation accuracy $\mathrm{Acc}(k)$ and the average token usage $\overline{T}(k)$ at each episode $k$. For comparison, we also include a Maj4 baseline: under the same truncated context, we sample four answers and take a majority vote, measuring $\mathrm{Maj@4}(k)$ and its token cost. Plotting “accuracy vs. episodes” and “token cost vs. episodes” reveals how model performance varies with question difficulty and compute budget, and highlights the relative merits of sequential decoding versus voting.
@@ F.2 Full Results of Effectiveness and Efficiency Analysis
To evaluate the proposed L2T, we conduct experiments on various benchmarks, including mathematical and code-related tasks, and across base models of different scales. In the main text, we report the performance of L2T on mathematical reasoning tasks of varying complexity. Table 1, Figure 2, and Figure 3 show that, compared to prior outcome-reward or process-reward methods, L2T delivers superior reasoning with less test-time compute. In this subsection, we assess its performance on more based models and more reasoning tasks, e.g., code generation tasks. Notably, mathematical reasoning and code generation serve as classic benchmarks for testing an LLM’s complex reasoning ability [43, 16, 32]. First, we report performance across additional model scales. The results for inference accuracy and compute efficiency are shown in Table 2 and Figure 7. We observe the same conclusion: L2T achieves state-of-the-art performance, attaining the highest inference accuracy with the smallest token budget. These findings demonstrate the broad effectiveness of our approach across models of varying scales. Secondly, we provide the performance of the proposed framework on the code generation task, and we evaluate the improvement of L2T on the LLM reasoning performance on the code generation task. Specifically, we run the GRPO/L2T fine-tuning pipeline on Qwen2-7B-Instruct and evaluate it using the standard HumanEval protocol. We set the temperature to 0.6 and top-p to 0.95 and generate 64 solutions per question, with a 1s timeout per attempt. We report the proportion of questions that pass all unit tests at least once. All tests run with a single-card batch size of 1 and a fixed random seed to ensure reproducibility. From the results in Figure 9, we can observe that compared with the outcome-reward-based RL method, L2T achieves better reasoning performance with less token budget. This proves the superiority of our proposed L2T and dense process rewards.
@@ F.3 Full Results of Ablation Studies
In Subsection 5.3, we conduct ablation studies to evaluate the optimal configuration and parameter settings. Since models in practical applications may face tasks of varying complexity, we further analyze these alternatives and assess batch size for a more comprehensive evaluation.
Considering the impact of batch size on model performance, particularly during fine-tuning based on the GRPO framework, we conduct an ablation study on batch size. Given the computational resource overhead and limits, we evaluate batch sizes of 64, 128, and 256 on A100 GPU clusters. The results, shown in Figure 9, indicate that a larger batch size can improve model performance to some extent. It is worth noting that while a larger batch size increases GPU memory usage, especially in RL-based optimization, our focus is on the token budget and accuracy during inference.
Considering questions of varying complexity, we perform evaluation on multiple benchmarks. We conduct an ablation study on the three core components of L2T, evaluating the contribution of each one by constructing alternative configurations: (i) Replacing information gain (config 1): replacing the proposed process reward with a task-aligned pre-trained reward model; (ii) Removing parameter compression penalty (config 2): completely removing the parameter compression penalty driven by the Fisher information matrix; (iii) Replacing low-rank approximation (config 3): using random sampling of 30% of network layers to approximate the Fisher information matrix, instead of the original low-rank approximation method. All ablation configurations follow the same hyperparameters and test protocols as the main experiment, each Tier is repeated five times to report average accuracy and average token consumption. The results in Figure 10 confirm similar conclusions as in the main text: replacing information gain with task-specific reward leads to an average accuracy drop of more than 3%, with a slight increase in token consumption; removing the parameter compression penalty results in about a 12% increase in consumption and a drop in accuracy; while random layer sampling reduces approximation overhead, the accuracy drops significantly, and the fluctuations increase substantially. These results validate the crucial role of each proposed component within L2T.

### v2

@@ 6 Conclusion
The appendix is organized as follows:
@@ Appendix A Proofs
Proof.
For a given state $s_{k}$, the posterior distribution $p(\tilde{\theta}_{k}\mid s_{k})\propto p(s_{k}\mid\tilde{\theta}_{k})\;p(\tilde{\theta}_{k})$, take the logarithm to get the unstandardized posterior $L(\tilde{\theta}_{k})=\log p(s_{k}\mid\tilde{\theta}_{k})+\log p(\tilde{\theta}_{k})$. Among them,
@@ Appendix B Pseudo-Code
[1] Input: Task distribution $P_{task}$; base model $\pi_{\theta}$;
@@
for iteration = $1,\cdots,T$ do
     Sample tasks $X\sim P_{task}$
     for each $x\in X$ do
         for each episode $k$ do
              Sample $N$ rollouts $\{z_{k}^{i},y_{k}^{i}\}\}_{i=1}^{N}\sim\pi_{\theta}(\cdot|s_{k})$
              Compute average progress reward $r_{k}^{\rm prg}$ via Definition 4.1 and Theorem 4.2
              Compute average outcome reward $r_{k}^{\rm out}$ following [16]
         end for
     end for
     Update the policy $\pi_{\theta}$ via Eq.5
end for
return solution
@@ Appendix D Benchmark Datasets
In this section, we briefly introduce all datasets used in our experiments. In summary, the benchmark datasets can be divided into two categories: (i) reasoning tasks for mathematical derivation, including AIME24-25, AMC, MATH500 [18], MinervaMATH [27], and Omni-MATH [13]; and (ii) reasoning tasks for code generation via HumanEval [8] The compositions of these benchmarks are as follows:
  • AIME24-25 comprises 30 questions from the 2024 and 2025 American Invitational Mathematics Examination, with 15 fill-in-the-blank questions per exam. These questions are more difficult than AMC, spanning number theory, combinatorics, geometry, and algebra.
@@
  • MinervaMATH comprises 12,500 high-school-level contest questions. Each includes detailed solution steps and spans prealgebra through precalculus.
@@ Appendix E Implementation Details
The implementation of our approach is built upon the openRL codebase and Llama-Factory codebase, with custom modifications to the loss function, i.e., replacing the reward function to our proposed reward mentioned in Subsection 4.2. For model training, we directly load the base models from Hugging Face, including DeepScaleR-1.5B-Preview, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeekR1-Distill-Qwen-7B, and Qwen2-7B-Instruct. For different reasoning tasks, we introduce the experimental settings in the corresponding sections of Section 5 and Appendix F. Unless otherwise specified, we follow the protocol of each benchmark and record the maj@4 results across different models. The training configuration is: the learning rate is set to $1.0\times 10^{-6}$, with a cosine learning rate scheduler, and a warm-up ratio of 0.1. We use a batch size of 256, with a maximum prompt length of 4,096 tokens and a maximum completion length of 16,384 tokens. The model is trained for 1 epoch, up to 10 epochs. Additionally, we set the ‘use_vllm’ flag to True to enable vLLM acceleration, with a GPU memory utilization of 0.8. We also utilize mixed precision training with BF16 enabled. The parameters for compression penalty approximation are handled by a single-layer MLP, with a Fisher information matrix damping factor set to $10^{-5}$. The regularization hyperparameters $\alpha$ and $\beta$ are set to 0.8 and 0.6 according to grid search results, respectively. Also, $\alpha$ can be set to 1 for simplicity, with the performance drop less than 1%. More evaluation of implementation is provided in Appendix F. The entire training is conducted on A100 GPU clusters, ensuring scalability and high computational efficiency. The main code is provided in supplementary materials.
@@ Appendix F Additional Experiments and Full Results
  +outcome-reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (-1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
  +length penalty 40.3 (-2.5) 30.3 (-6.4) 77.3 (-5.7) 83.2 (-2.0) 23.0 (-1.6) 50.8 (-3.7)
@@ F.1 More Details and Results of the Motivating Experiments
In Subsection 3.2, we evaluate how efficiently existing methods use tokens. We benchmark two base models, DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B, on Omni-MATH (4,428 questions across 33+ subfields, split into Tiers 1-4 by expert difficulty labels). Both models have been fine-tuned with outcome-based RL. To study performance at varying reasoning depths, we split each generated reasoning chain into up to $K=20$ episodes using ‘<think>…</think>’, then record the sequential-generation accuracy $\mathrm{Acc}(k)$ and the average token usage $\overline{T}(k)$ at each episode $k$. For comparison, we also include a Maj4 baseline: under the same truncated context, we sample four answers and take a majority vote, measuring $\mathrm{Maj@4}(k)$ and its token cost. Plotting “accuracy vs. episodes” and “token cost vs. episodes” reveals how model performance varies with question difficulty and compute budget, and highlights the relative merits of sequential decoding versus voting.
@@ F.2 Full Results of Effectiveness and Efficiency Analysis
To evaluate the proposed L2T, we conduct experiments on various benchmarks, including mathematical and code-related tasks, and across base models of different scales. In the main text, we report the performance of L2T on mathematical reasoning tasks of varying complexity. Table 1, Figure 2, and Figure 3 show that, compared to prior outcome-reward or process-reward methods, L2T delivers superior reasoning with less test-time compute. In this subsection, we assess its performance on more based models and more reasoning tasks, e.g., code generation tasks. Notably, mathematical reasoning and code generation serve as classic benchmarks for testing an LLM’s complex reasoning ability [43, 16, 32]. First, we report performance across additional model scales. The results for inference accuracy and compute efficiency are shown in Table 2 and Figure 7. We observe the same conclusion: L2T achieves state-of-the-art performance, attaining the highest inference accuracy with the smallest token budget. These findings demonstrate the broad effectiveness of our approach across models of varying scales. Secondly, we provide the performance of the proposed framework on the code generation task, and we evaluate the improvement of L2T on the LLM reasoning performance on the code generation task. Specifically, we run the GRPO/L2T fine-tuning pipeline on Qwen2-7B-Instruct and evaluate it using the standard HumanEval protocol. We set the temperature to 0.6 and top-p to 0.95 and generate 64 solutions per question, with a 1s timeout per attempt. We report the proportion of questions that pass all unit tests at least once. From the results in Figure 8, we can observe that compared with the outcome-reward-based RL method, L2T achieves better reasoning performance with less token budget. This proves the superiority of our proposed L2T and dense process rewards.
@@ F.3 Full Results of Ablation Studies
In Subsection 5.3, we conduct ablation studies to evaluate the optimal configuration and parameter settings. Considering questions of varying complexity, we perform evaluation on multiple benchmarks. We conduct an ablation study on the three core components of L2T, evaluating the contribution of each one by constructing alternative configurations: (i) Replacing information gain (config 1): replacing the proposed process reward with a task-aligned pre-trained reward model; (ii) Removing parameter compression penalty (config 2): completely removing the parameter compression penalty driven by the Fisher information matrix; (iii) Replacing low-rank approximation (config 3): using random sampling of 30% of network layers to approximate the Fisher information matrix, instead of the original low-rank approximation method. All ablation configurations follow the same hyperparameters and test protocols as the main experiment, each Tier is repeated five times to report average accuracy and average token consumption. The results in Figure 9 confirm similar conclusions as in the main text: replacing information gain with task-specific reward leads to an average accuracy drop of about 1.9%, with a slight increase in token consumption; removing the parameter compression penalty results in about a 12% increase in consumption and a drop in accuracy; while random layer sampling reduces approximation overhead, the accuracy drops significantly, and the fluctuations increase substantially. These results validate the crucial role of each proposed component within L2T.